Circular No. 6847

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

                URL http://cfa-www.harvard.edu/iau/cbat.html

Phone 617-495-7244/7440/7444 (for emergency use only)

SUPERNOVA 1998X IN NGC 6754 Simon Woodings, Ralph Martin, and Andrew Williams report their dis-

covery of an apparent supernova (R = 17) on a CCD frame taken on Mar. 13.88 UT (confirmed on Mar. 15.86) as part of the Perth Astronomy Research Group’s automated supernova search with the 0.61-m Perth-Lowell telescope. SN 1998X is located 23′′.8 ± 1′′ west and 3′′.1 ± 1′′ south of the center of NGC 6754 (α = 19^{h}11^{m}.4, δ = −50^{o}39′, equinox 2000.0).

F. Patat, European Southern Observatory, (ESO); and M. Della Valle, Padova University, write: “We have observed SN 1998X with the ESO 1.5-m telescope (+ Boller & Chivens spectrograph; exposure time 45 min; range 320–710 nm, resolution 0.5 nm) on Mar. 23.40 UT at La Silla. A preliminary reduction of the CCD spectrogram shows that this object is a type-II supernova, ∼ 3–4 months after explosion. The spectrum consists of a multitude of lines with the usual P-Cyg profiles, superimposed on a red continuum. Most of the lines arise from the H I Balmer series and Na I, Sc II, and Fe II. The expansion velocities deduced from the minima of the most prominent features are 6100 (Hα), 3050 (Na I), and 3570 km/s (Fe II). The recession velocity at the supernova location is 3110 km/s, as deduced from the narrow Hα emissions. A narrow interstellar absorption of Na I D, arising in the parent galaxy, is clearly visible at 595.3 nm.”

NOVA SAGITTARII 1998 S. Nakano, Sumoto, Japan, forwards the following position and magni-

tude for this object (cf. IAUC 6846) from Y. Kushida, Yatsugatake South Base Observatory (0.40-m f/6 reflector), from an image obtained on Mar. 23.771 UT: α = 18^{h}21^{m}40^{s}.47, δ = −27^{o}31′38′′.0 (equinox 2000.0), V = 8.1. Nakano also reports the following position end figures from M. Aoki, Toyama (0.43-m f/6 reflector), who reported V = 7.8 on Mar. 23.724: 40^{s}.47, 37′′.9. The closest star present in the USNO A1.0 catalogue (and on the Digital Sky Survey) is one of mag ∼ 17 located 6′′.5 north of the nova.

G. J. Garradd, Loomberah, N.S.W., reports the following position end figures and magnitude from an unfiltered CCD image obtained on Mar. 23.682 UT with a 0.25-m f/4.1 reflector: 40^{s}.51, 38′′.5, 8.2.

SUPERNOVA 1998S IN NGC 3877 Revised position end figures and magnitude from J.-C. Merlin, regard-

ing line 11 of IAUC 6831: 06^{s}.11, 55′′.8, 12.5.

1998 March 23  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1998 CBAT  Daniel W. E. Green